# Reference.bib

### v1

@@ 4.2. Definition of CMCX
A combinatorial monodromy complex (CMCX) $\mathcal{T}$ associated to $\mathbb{V}\rightarrow B$ and $\hat{B}$ is defined by assigning each $I\in\mathcal{I}$ with $Z_{I}\neq\emptyset$ a polyhedral complex $\Sigma_{\mathcal{T}}(I)$ supported on the closed combinatorial monodromy cone $\overline{\boldsymbol{\sigma}_{I}}\subset\mathbb{Q}^{|I|}/\mathrm{Ker}(\lambda_{I})$, such that for any $I^{\prime}\leq I$, the adjacent map $\xi_{I,I^{\prime}}$ induces an embedding of subcomplexes $\Sigma_{\mathcal{T}}(I^{\prime})\leq\Sigma_{\mathcal{T}}(I)$.
@@
We say the completion $\hat{B}$ of $B$ is simplicial corresponding to the $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow B$ or just simplicial, if for any $b\in Z_{I}^{*}\subset\hat{B}$ the CMC $\boldsymbol{\sigma}_{b}$ is a simplicial cone, and for any $i\in I$ the local monodromy logarithm $N_{i}$ (view as a CMC) is a face of $\boldsymbol{\sigma}_{b}$ via the adjacent map $\xi_{I,i}$.
@@
If $\hat{B}$ is simplicial corresponds to $\mathbb{V}\rightarrow B$, then the collection of cones $\mathcal{T}_{0}$ associating each strata $Z_{I}^{*}\subset\hat{B}-B$ the polyhedral complex {All faces of $\{\text{All faces of }\sigma_{I}\}$ is a CMCX. In particular, this includes the case when $B$ is a curve or a surface.
@@
In these cases it is straight-forward to verify that every adjacent map $\xi_{I,I^{\prime}}$ induces an embedding of sub-complexes. ∎
@@ 4.3. CMCX and logarithmic modifications
We call the map $\eta_{I,J}$ (or $\eta_{b,b^{\prime}}$) as the descendant map.
With the definition of the descendant map, one may define the compatibility of $\mathcal{T}$ and the logarithmic modification $\hat{B}\rightarrow B$ as follows.
@@
We say a CMCX $\mathcal{T}$ associated to $\overline{B}$ is compatible with the logarithmic modification $\hat{B}\rightarrow\hat{B}$ (or vise versa) if for any $b^{\prime}\in\overline{B}$ lying above $b\in\hat{B}$, there exists a unique $\boldsymbol{\tau}_{b}\in\Sigma_{\mathcal{T}}(b)$ [^3] For convention, by $\Sigma_{\mathcal{T}}(b)$ we mean $\Sigma_{\mathcal{T}}(I)$ for the strata $I$ containing $b$. such that $\eta_{b^{\prime},b}(\boldsymbol{\sigma}_{b^{\prime}})\subset\boldsymbol{\tau}_{b}$. Moreover, this unique $\boldsymbol{\tau}_{b}$ satisfies $(\tau_{b},F_{b^{\prime}})$ is a nilpotent orbit.
@@
Starting from this point, When a descendant map $\eta_{I,J}$ or an adjacent map $\xi_{I,J}$ are clear, we ignore the map and denote the corresponding map as $\boldsymbol{\sigma}_{J}\hookrightarrow\overline{\boldsymbol{\sigma}_{I}}$ or $\boldsymbol{\sigma}_{J}\subset\overline{\boldsymbol{\sigma}_{I}}$.
If we assume the global monodromy group $\Gamma$ is abelian, a CMCX is equivalent to a $\Gamma$-compatible fan $\Sigma$ given by [KU08, Theorem 4.3.1(iii)]. In general, CMCX is different from the notation of fan as it does not require compatibility with global monodromy.
Suppose a CMCX $\mathcal{T}$ corresponds to the $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow B$ and the completion $B\subset\hat{B}$ is given. The following theorem is the first technical result of this paper. The proof requires backgrounds in logarithmic geometry or toroidal embedding. References include [KU08] and [KKMSD89].
@@
There exists a logarithmic modification $(\overline{B},B)$ of $(\hat{B},B)$ compatible with $\mathcal{T}$.
@@
Let $\boldsymbol{C}_{\hat{B}}$ be the Kato fan (or toroidal fan) associated to the toroidal embedding $(\hat{B},B)$. It associates each $I\in\mathcal{I}$ a (closed) simplicial cone $\boldsymbol{C}_{I}$ with dimension same as the codimension of $Z_{I}^{*}\subset\hat{B}$ because of the SNC property. There is a combinatorial monodromy map $\mathfrak{m}_{I}:\boldsymbol{C}_{I}\rightarrow\overline{\boldsymbol{\sigma}_{I}}$ sending the $1$-dimensional face of $\boldsymbol{C}_{I}$ corresponding to the divisor $Z_{i}$ to the monodromy logarithm $\boldsymbol{N}_{i}\subset\overline{\boldsymbol{\sigma}_{I}}$ (viewed as a CMC), and extended linearly.
For each $I\in\mathcal{I}$, since $\Sigma_{\mathcal{T}}(I)$ is a polyhedral decomposition of the CMC $\boldsymbol{\sigma}_{I}$, its pre-image under $\mathfrak{m}_{I}$ gives a polyhedral decomposition of $\boldsymbol{C}_{I}$ which we denote as $\boldsymbol{C}^{\prime}_{I}$. This gives a logarithmic modification $\hat{B}(I)\rightarrow\hat{B}$.
Consider the collection of logarithmic modifications $\{\hat{B}(I)\rightarrow\hat{B}\}_{I\subset\mathcal{I}}$. For any $I,J\subset\mathcal{I}$ with $Z_{I\cup J}\neq\emptyset$, Both $\hat{B}(I)\rightarrow\hat{B}$ and $\hat{B}(J)\rightarrow\hat{B}$ factor through $\hat{B}(I\cup J)\rightarrow\hat{B}$ because of the property of CMCX. Therefore, the collection of logarithmic modifications $\{\hat{B}(I)\rightarrow\hat{B}\}_{I\subset\mathcal{I}}$ agree on the overlaps. It is a classical result in logarithmic or toroidal geometry that there exists a logarithmic modification $\overline{B}\rightarrow\hat{B}$ which factors through every $\hat{B}(I)\rightarrow\hat{B}$, and by refining the Kato fan if necessary, we may assume $\overline{B}$ is smooth with $\overline{B}-B$ an SNC divisor.
Finally, for any $b^{\prime}\in Z_{J}^{*}\subset\overline{B}$ lying above $b\in Z_{I}^{*}\subset\hat{B}$, by the construction of $\hat{B}$ there exists a unique cone $\boldsymbol{c}_{I}\in\boldsymbol{C}^{\prime}_{I}$ such that locally around $b\in\hat{B}$, $b^{\prime}$ lies above the boundary component of $\hat{B}(I)\rightarrow\hat{B}$ determined by $\boldsymbol{c}_{I}$. The cone $\boldsymbol{c}_{I}$ is mapped to a CMC $\boldsymbol{\tau}_{b^{\prime}}\in\Sigma_{\mathcal{T}}(I)$, and it is clear that $(\tau_{b^{\prime}},F_{b^{\prime}})$ must be a nilpotent orbit as $(\sigma_{I},F_{b^{\prime}})$ is. ∎
@@ 5. Local monodromy and proper Kato–Nakayama–Usui maps
During this section we fix a logarithmic modification $\overline{B}\rightarrow\hat{B}$ as well as a CMCX $\mathcal{T}$ compatible with $\overline{B}$. The main goal of this section is to show the following local extension result. Write $\overline{B}-B=\bigcup_{J\subset\overline{\mathcal{I}}}Z_{J}$ as the union of its smooth boundary strata, and $Z_{J}^{*}:=Z_{J}-\cup_{J\not\subset K}Z_{K}$.
Theorem 5.1.
There exists a finite open cover $\overline{B}=\bigcup_{k}W_{k}$ satisfying the following properties.
  1. (1)
    For each $k$ the monodromy group $\Gamma_{k}$ of $\mathbb{V}\rightarrow B$ restricted to $U_{k}:=W_{k}\cap B$ is abelian.
  2. (2)
    For each $U_{k}$, there exists a KNU extension
    (5.1) ${U_{k}}$${\Gamma_{k}\backslash D}$${W_{k}}$${\Gamma_{k}\backslash D_{\tau_{k}}}$$\scriptstyle{\Phi_{k}}$$\scriptstyle{\overline{\Phi}_{k}}$
    such that each $\overline{\Phi}_{k}$ is proper. In particular, $W_{k}$ may be chosen as a union of compact fibers of $\overline{\Phi}_{k}$.
  3. (3)
    For each $\overline{\Phi}_{k}:W_{k}\rightarrow\Gamma_{k}\backslash D_{\tau_{k}}$ and $b^{\prime}\in W_{k}$ lying above $b\in\hat{B}$, up to some finite data, the image $\overline{\Phi}_{k}(b^{\prime})$ records the nilpotent orbit $(\tau_{b^{\prime}},F_{b^{\prime}})$ where $\tau_{b^{\prime}}$ is isomorphic to some $\boldsymbol{\tau}_{b^{\prime}}\in\Sigma_{\mathcal{T}}(b)$.
  4. (4)
    Suppose for $k,l$, $F_{k}$ and $F_{l}$ are fibers of $\overline{\Phi}_{k}$ and $\overline{\Phi}_{l}$ correspondingly such that $F_{k}\cap F_{l}\neq\emptyset$, then $F_{k}\cap F_{l}$ is a union of connected components of both $F_{k}$ and $F_{l}$.
The remaining of this section will be devoted to the construction of an open cover $\overline{B}=\bigcup_{k}W_{k}$ with all of the properties required by Theorem 5.1.
@@ 5.1. The Nilpotent Cone Closure (NCC)
Let $J\subset\overline{\mathcal{I}}$ be a subset such that $Z_{J}\neq\emptyset$. For any point $y\in Z_{J}^{*}$ lying above some $b\in Z_{I}^{*}\subset\hat{B}$, by Definition 4.10 there exists a unique $\boldsymbol{\tau}_{J}\in\Sigma_{\mathcal{T}}(I)$ such that $\boldsymbol{\sigma}_{y}=\boldsymbol{\sigma}_{J}\subset\boldsymbol{\tau}_{J}$.
@@
For any $J\subset\overline{\mathcal{I}}$, the nilpotent cone closure (NCC) of $J$ relative to $\mathcal{T}$ or simply the NCC of $J$, denoted as $\mathfrak{C}_{\mathcal{T}}(J)$ or simply just $\mathfrak{C}(J)$ when $\mathcal{T}$ is fixed, is the unique maximal collection of subsets of $\overline{\mathcal{I}}$ such that:
  1. (1)
    $J\in\mathfrak{C}(J)$.
  2. (2)
    $\bigcup_{K\in\mathfrak{C}(J)}Z_{K}^{*}$ is connected.
  3. (3)
    For any $K\in\mathfrak{C}(J)$, there exists $K^{\prime}\in\mathfrak{C}(J)$ with $K\sim K^{\prime}$.
@@
We present an algorithm for determining $\mathfrak{C}(J)$ with a given $J\subset\overline{\mathcal{I}}$.
Lemma 5.6.
For a given $J\subset\overline{\mathcal{I}}$, the following algorithm determines the set $\mathfrak{C}(J)$.
  1. Step 1:
    Starting with the set $\mathfrak{C}(J)_{0}:=\{J\}$.
  2. Step 2:
    Given the set $\mathfrak{C}(J)_{k}$ with $k\geq 0$, construct the set $\mathfrak{C}(J)_{k+1}:=\{K\subset\overline{\mathcal{I}},K\sim K^{\prime}\ \text{for some }K^{\prime}\in C(J_{k})\}$.
  3. Step 3:
    If $\mathfrak{C}(J)_{m}=\mathfrak{C}(J)_{m+1}$, terminate the algorithm and output $\mathfrak{C}(J):=\mathfrak{C}(J)_{m}$. Otherwise, back to Step 2.
Proof.
Since $\overline{\mathcal{I}}$ is finite, the algorithm must terminate in finitely many steps. By Definition 5.5, the output is exactly the set $\mathfrak{C}(J)$. ∎
@@ 5.2. Connected Limiting Fiber (CLF)
Fix some $J\subset\overline{\mathcal{I}}$ such that $Z_{J}\neq\emptyset$. For any $p\in Z_{J}^{*}$, we may choose a coordinate neighborhood $p\in B_{p}\subset\overline{B}$ such that $B_{p}\cap B$ is homeomorphic to a product of (punctured and unpunctured) disks. Let $\mathcal{U}_{p}^{-}:=B_{p}\cap(B\cup Z_{I}^{*})$ and $\mathcal{U}_{p}:=\mathcal{U}_{p}^{-}\cap B$, we have the local KNU map $\Phi_{p}^{\mathrm{KNU}}$ extending the local period map $\Phi_{p}$:
(5.2) ${\mathcal{U}_{p}}$${\Gamma_{\sigma_{J}}\backslash D}$${\Gamma_{\tau_{J}}\backslash D}$${\mathcal{U}_{p}^{-}}$${\Gamma_{\sigma_{J}}\backslash D_{\sigma_{J}}}$${\Gamma_{\tau_{J}}\backslash D_{\tau_{J}}}$$\scriptstyle{\Phi_{p}}$$\scriptstyle{\Phi_{p}^{\mathrm{KNU}}}$
@@
Let $A\subset\overline{B}-B$ be a closed subvariety. We say $A$ is a connected limiting fiber (CLF) if the following conditions hold:
  1. (1)
    $A$ is connected.
  2. (2)
    For any $J\subset\overline{\mathcal{I}}$ such that $A\cap Z_{J}^{*}\neq\emptyset$, any $p\in A\cap Z_{J}^{*}$ and any coordinate neighborhood $p\in B_{p}\subset\overline{B}$ over which the diagram (5.2) exists, $A\cap\mathcal{U}_{p}^{-}$ is a $\Phi_{p}^{\mathrm{KNU}}$-fiber.
In this case, we denote $A_{J}:=A\cap Z_{J}^{*}$ as the $J$-component of $A$.
Now we fix some $J\subset\mathcal{I}$ and some CLF $A$ with $A_{J}\neq\emptyset$. We will show:

### v4

@@ 4.2. Definition of CMCX
A combinatorial monodromy complex (CMCX) $\mathcal{T}$ associated to $\mathbb{V}\rightarrow B$ and $\hat{B}$ is defined by assigning each $I\in\mathcal{I}$ with $Z_{I}\neq\emptyset$ a polyhedral complex $\Sigma_{\mathcal{T}}(I)$ supported on the closed combinatorial monodromy cone $\boldsymbol{\sigma}_{I}\subset\mathbb{Q}^{|I|}/\mathrm{Ker}(\lambda_{I})$, such that for any $I^{\prime}\leq I$, the adjacency map $\xi_{I,I^{\prime}}$ induces an embedding of subcomplexes $\Sigma_{\mathcal{T}}(I^{\prime})\leq\Sigma_{\mathcal{T}}(I)$.
@@
We say the completion $\hat{B}$ of $B$ is simplicial corresponding to the $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow B$ or just simplicial, if for any $b\in Z_{I}^{*}\subset\hat{B}$ the CMC $\boldsymbol{\sigma}_{b}$ is a simplicial cone, and for any $i\in I$ the local monodromy logarithm $N_{i}$ (view as a CMC) is a face of $\boldsymbol{\sigma}_{b}$ via the adjacency map $\xi_{I,i}$. We say $\mathbb{V}\rightarrow B$ has simplicial local monodromy or just say $\mathbb{V}\rightarrow B$ is simplicial if such a completion $\hat{B}$ of $B$ exists.
@@
If $\hat{B}$ is simplicial corresponds to $\mathbb{V}\rightarrow B$, then the collection of cones $\mathcal{T}_{0}$ associating each strata $Z_{I}^{*}\subset\hat{B}-B$ the polyhedral complex $\{\text{All faces of }\sigma_{I}\}$ is a CMCX. In particular, this includes the case when $B$ is a curve or a surface.
@@
In these cases it is straight-forward to verify that every adjacency map $\xi_{I,I^{\prime}}$ induces an embedding of sub-complexes. ∎
@@ 4.3. CMCX and logarithmic modifications
We refer to the map $\eta_{I,J}$ (or $\eta_{b,b^{\prime}}$) as the descendant map.
With the definition of the descendant map, one may define the compatibility of $\mathcal{T}$ and the logarithmic modification $\overline{B}\rightarrow\hat{B}$ as follows.
@@
We say a CMCX $\mathcal{T}$ associated to $\overline{B}$ is compatible with the logarithmic modification $\overline{B}\rightarrow\hat{B}$ (or vice versa) if for any $b^{\prime}\in\overline{B}$ lying above $b\in\hat{B}$, there exists a unique $\boldsymbol{\tau}_{b}\in\Sigma_{\mathcal{T}}(b)$ [^3] For convention, by $\Sigma_{\mathcal{T}}(b)$ we mean $\Sigma_{\mathcal{T}}(I)$ for the strata $I$ containing $b$. such that $\eta_{b^{\prime},b}(\boldsymbol{\sigma}_{b^{\prime}}^{\circ})\subset\boldsymbol{\tau}_{b}^{\circ}$. Moreover, this unique $\boldsymbol{\tau}_{b}$ satisfies $(\tau_{b},F_{b^{\prime}})$ is a nilpotent orbit.
@@
Starting from this point, When a descendant map $\eta_{I,J}$ or an adjacency map $\xi_{I,J}$ are clear, we ignore the map and denote the corresponding inclusions as $\boldsymbol{\sigma}_{J}\hookrightarrow\boldsymbol{\sigma}_{I}$ or $\boldsymbol{\sigma}_{J}\subset\boldsymbol{\sigma}_{I}$.
If we assume the global monodromy group $\Gamma$ is abelian, a CMCX is equivalent to a $\Gamma$-compatible fan $\Sigma$ given by [KU08, Theorem 4.3.1(iii)]. In general, the notion of a CMCX is different from the notation of fan as it depends more on the base of the map, and in particular does not require any kind of compatibility with global monodromy.
We breifly recall some of the literature of toroidal embeddings and toroidal/logarithmic modifications. Suppose a CMCX $\mathcal{T}$ corresponds to the $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow B$ and the completion $B\subset\hat{B}$ with $\hat{B}-B$ an SNC divisor are given.
In the theory of [Kem+73], the pair $(\hat{B},B)$ is a toroidal embedding. In particular, there exists a canonically-constructed toroidal fan $\boldsymbol{C}$ associated to the toroidal embedding as [Kem+73, pg. 71] which, in our case, is a topological polyhedral complex consisting of simplicial cones.
If $\boldsymbol{C}^{\prime}$ is another complex refining $\boldsymbol{C}$, then there is another toroidal embedding pair $(\hat{B}^{\prime},B)$ with toroidal fan $\boldsymbol{C}^{\prime}$ and a proper birational morphism $\hat{B}^{\prime}\rightarrow\hat{B}$ restricting to the identity map on $B$. This is called a toroidal modification. As a summary, similar to the case for toric varieties, we have the correspondence:
{Finite “allowable” subdivisions of toroidal fans $\Leftrightarrow$ {Toroidal modifications},
see [Kem+73, pg. 90]. In our case, “allowable” subdivisions are exactly finite rational polyhedral subdivisions.
The classical theory of toroidal embeddings and modifications gneeralize to the logarithmic setting, see [KU08, Chap. 3.6] for an overview. In particular, the term “logarithmic modification” has the same meaning as “toroidal modification” in our case, and we will use them interchangably.
With the general philosophy of logarithmic modifications and polyhedral subdivision of toroidal fans, we are ready to prove the first technical result of this paper below. Note that the proof uses definitions of semicomplex and the canonical complexification which are not defined until Section 8.
@@
Given a CMCX $\mathcal{T}$ on $\hat{B}$, there exists a logarithmic modification $(\overline{B},B)$ of $(\hat{B},B)$ compatible with $\mathcal{T}$.
@@
Let $\boldsymbol{C}_{\hat{B}}$ be the toroidal fan associated to the toroidal embedding $(\hat{B},B)$. It associates to each $I\in\mathcal{I}$ a closed simplicial cone $\boldsymbol{C}_{I}$ with dimension same as the codimension of $Z_{I}^{*}\subset\hat{B}$ because of the SNC property. There is a combinatorial monodromy map $\mathfrak{m}_{I}:\boldsymbol{C}_{I}\rightarrow\boldsymbol{\sigma}_{I}$ sending the $1$-dimensional face of $\boldsymbol{C}_{I}$ corresponding to the divisor $Z_{i}$ to the monodromy logarithm $\boldsymbol{N}_{i}\subset\boldsymbol{\sigma}_{I}$ (viewed as a CMC), and extended linearly.
For each $I\in\mathcal{I}$, since $\Sigma_{\mathcal{T}}(I)$ is a polyhedral decomposition of the CMC $\boldsymbol{\sigma}_{I}$, its pre-image under $\mathfrak{m}_{I}$ gives a polyhedral decomposition of $\boldsymbol{C}_{I}$ which we denote as $\boldsymbol{C}_{\mathcal{T}}(I)$. By Definition 8.8 and Lemma 8.9, the collection obtained from $\boldsymbol{C}_{\hat{B}}$ by replacing every $\boldsymbol{C}_{K}\leq\boldsymbol{C}_{I}$ with the set $\boldsymbol{C}_{\mathcal{T}}(I)|_{\boldsymbol{C}_{K}}$ is a semicomplex. Denote its canonical complexification by Definition 8.11 as $\boldsymbol{C}_{I}^{\prime}$. Since $\boldsymbol{C}_{I}^{\prime}$ is a refinement of $\boldsymbol{C}_{I}$, it gives a logarithmic modification $\hat{B}(I)\rightarrow\hat{B}$. Let $\boldsymbol{C}_{\overline{B}}$ be a common refinement of all $\boldsymbol{C}_{I}^{\prime}$ and $\overline{B}\rightarrow\hat{B}$ be the corresponding logarithmic modification. By further resolving $\overline{B}$ if necessary, we may assume $\overline{B}$ is smooth and $\overline{B}-B$ is an SNC divisor.
We show the logarithmic modification $\overline{B}\rightarrow\hat{B}$ has the required properties. For any $b^{\prime}\in Z_{J}^{*}\subset\overline{B}$ lying above $b\in Z_{I}^{*}\subset\hat{B}$, by the construction of $\hat{B}$ there exists a unique cone $\boldsymbol{c}_{I}\in\boldsymbol{C}^{\prime}_{I}$ such that locally around $\in\hat{B}, bb^{\prime}$ lies above the boundary component of $\hat{B}(I)\rightarrow\hat{B}$ determined by $\boldsymbol{c}_{I}$. The cone $\boldsymbol{c}_{I}$ is mapped into a CMC $\boldsymbol{\tau}_{b^{\prime}}\in\Sigma_{\mathcal{T}}(I)$, and it is clear that $(\tau_{b^{\prime}},F_{b^{\prime}})$ must be a nilpotent orbit as $(\sigma_{I},F_{b^{\prime}})$ is. ∎
@@ 5. Local monodromy and proper Kato–Nakayama–Usui maps
During this section we fix a logarithmic modification $\overline{B}\rightarrow\hat{B}$ as well as a CMCX $\mathcal{T}$ compatible with $\overline{B}$. We define $\overline{\mathcal{I}}$ is the collection of boundary strata of $\overline{B}$. Write $\overline{B}-B=\bigcup_{J\subset\overline{\mathcal{I}}}Z_{J}$ as the union of its smooth boundary strata, and $Z_{J}^{*}:=Z_{J}-\cup_{J\not\subset K}Z_{K}$.
@@ 5.1. The Nilpotent Cone Closure (NCC)
Let $J\subset\overline{\mathcal{I}}$ be a subset such that $Z_{J}\neq\emptyset$. For any point $y\in Z_{J}^{*}$ lying above some $b\in Z_{I}^{*}\subset\hat{B}$, by Definition 4.10 there exists a unique $\boldsymbol{\tau}_{J}\in\Sigma_{\mathcal{T}}(I)$ such that $\boldsymbol{\sigma}_{y}=\boldsymbol{\sigma}_{J}\subset\boldsymbol{\tau}_{J}^{\circ}$.
@@
For any $J\subset\overline{\mathcal{I}}$, the nilpotent cone closure (NCC) of $J$ relative to $\mathcal{T}$, denoted as $\mathfrak{C}_{\mathcal{T}}(J)$ or simply just $\mathfrak{C}(J)$ when $\mathcal{T}$ is fixed, is the equivalence class of $J$ for the equivalence relation generated by $\sim$.
@@ 5.2. Connected Limiting Fiber (CLF)
For any index $K\subset\overline{\mathcal{I}}$ and $p\in Z_{K}^{*}$, we may choose a neighborhood $p\in B_{p}\subset\overline{B}$ such that $B_{p}\cap B$ is homeomorphic to a product of (punctured and unpunctured) disks. Let $\mathcal{U}_{p}:=\mathcal{U}_{p}^{-}\cap B$. Fix some $J\subset\overline{\mathcal{I}}$. Let $\mathcal{U}_{\mathfrak{C}(J)}\subset B$ be the punctured neighborhood of $Z_{\mathfrak{C}(J)}$ obtained by
$\mathcal{U}_{\mathfrak{C}(J)}:=\bigcup_{p\in Z_{\mathfrak{C}(J)}}\mathcal{U}_{p},\ \ B_{\mathfrak{C}(J)}:=\bigcup_{p\in Z_{\mathfrak{C}(J)}}B_{p}$
where for each $p$, $\mathcal{U}_{p}$ is a neighborhood supporting the existence of diagram (5.1). Moreover, by shrinking the $\mathcal{U}_{p}$ we may insist that $B_{\mathfrak{C}(J)}$ deformation retracts to $Z_{\mathfrak{C}(J)}$.
Let $\Gamma_{\mathfrak{C}(J)}$ be the monodromy group of $\mathbb{V}|_{\mathcal{U}_{\mathfrak{C}(J)}}\rightarrow\mathcal{U}_{\mathfrak{C}(J)}$.
Lemma 5.6.
$\Gamma_{\mathfrak{C}(J)}$ stabilizes $\tau_{J}\subset\mathfrak{g}_{\mathbb{Q}}$, and centralizes $\tau_{J}\subset\mathfrak{g}_{\mathbb{Q}}$ with neat $\Gamma$.
Proof.
Let $p\in Z_{K}^{*}$ for some $K\in\mathfrak{C}(J)$, and let $q\in U_{p}$. By construction we may assign to such a $q$ a cone $\tau_{q}\subset\mathrm{End}(\mathbb{V}_{q})$ corresponding to the monodromy cone around $p$, and its image in $\hat{B}$. Moreover, by the definition of the $\sim$ relation this assignment glues up to a sub-sheaf of monoids of $\mathrm{End}(\mathbb{V})\mid_{U_{\mathfrak{C}(J)}}$. Hence, we learn that the monodromy $\Gamma_{\mathfrak{C}(J)}$ stabilizes $\tau_{J}$.
For the final claim, note that neat groups are torsion-free. Since automorphisms of polyhedral cones are finite, the claim follows. ∎
Corollary 5.7.
There are well defined $KNU$ maps
[…]
Proof.
This is an immediate application of Corollary 3.11. Indeed, the fan consisting of $\tau$ and its faces is stabilized by $\Gamma_{\mathfrak{C}(J)}$ by Lemma 5.6, so we get a weak fan strongly compaible with $\Gamma_{\mathfrak{C}(J)}$. Moreover the definition of CMCX implies that the variation $\mathbb{V}$ restricted to $U_{\mathfrak{C}(J)}$ is compatible in the sense of Definition 3.9. Indeed, each relevant $\sigma^{\circ}$ is contained in a unique open $\tau^{\circ}$, and this induces a nilpotent orbit. Hence the corollary applies and the claim is proven. ∎
@@
Let $A\subset\overline{B}$ be an analytic subvariety. We say $A$ is a connected limiting fiber (CLF) if for some fixed $J\subset\overline{\mathcal{I}}$ and its associated $\Phi_{\mathfrak{C}(J)}^{\mathrm{KNU}}$, $A\subset Z_{\mathfrak{C}(J)}$ is a connected component of a fiber of $\Phi_{\mathfrak{C}(J)}^{\mathrm{KNU}}$.
Lemma 5.9.
The CLFs are compact.
Proof.
[…]
Let $\Gamma^{1}_{J}\subset\mathbf{G}(\mathbb{Z})$ be the arithmetic and centralizing $\sigma_{J}$ (and therefore preserving $W(\sigma_{J})$). Note that $\Gamma_{\mathfrak{C}(J)}\subset\Gamma^{1}_{J}$. By [Bak+24, Thm 6.4] there is a natural $\mathbb{R}_{\mathrm{alg}}$-definable structure on $\Gamma^{1}_{J}\backslash A(\sigma_{J})$ such that $\Phi_{J}:Z_{\mathfrak{C}(J)}\to\Gamma^{1}_{J}\backslash A(\sigma)$ is definable. Now $A$ is a connected component of a fiber of $\Phi_{J}$, and therefore $A$ must be $\mathbb{R}_{\mathrm{an},\exp}$-definable.
As a consequence $A\cap B_{q}$ has finitely many connected components, and let $A_{0}$ be one that contains $q$ in its closure.
[…]
5.3. Construction of the analytic structure
We define $\overline{\wp}$ to be the space of CLFs. There is a natural topology on $\overline{\wp}$ induced by the quotient map $\overline{\Phi}:\overline{B}\to\overline{\wp}$, which is trivially an extension of $\Phi$.
Our goal this subsection is to prove the following theorem:
Theorem 5.10.
The ringed space $(\overline{\wp},\overline{\Phi}_{*}\mathcal{O}^{\mathrm{an}}_{\overline{B}})$ is a normal, compact, holomorphic variety.
The rest of the section is devoted to the proof of Theorem 5.10, which clearly implies Theorem 2.8.
Proof.
First, we construct a suitable partition into subsets foliated by CLFs:
Lemma 5.11.
For each CLF $A\subset Z_{\mathfrak{C}(J)}$, there exists an open set $A\subset U\subset\mathcal{U}_{\mathfrak{C}(J)}$ such that $\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)}\!\!\mid_{U}:U\to\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)}(U)$ is a proper map. In particular, $U$ is foliated by CLFs.
Proof.
Let $w=\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)}(A)$. First, since $A$ is compact we may and do shrink $\mathcal{U}_{A}$ so that $A=(\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)})^{-1}(w)$.
[…]
Next, $w\in W$ be an open neighborhood such that $\overline{W}\cap W^{\prime}=\emptyset$. Then
$(\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)})^{-1}(\overline{W})\cap\overline{U^{\prime}}=(\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)})^{-1}(\overline{W})\cap U^{\prime}.$
We now set $U:=(\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)})^{-1}(\overline{W})\cap U^{\prime}$ and claim that this satisfies the assumption.
Indeed, the map $\overline{U^{\prime}}\to\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)}(\overline{U^{\prime}})$ is a proper map, and $\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)}\!\!\mid_{U}:U\to\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)}(U)$ is simply the base change under the inclusion $\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)}(U)\to\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)}(\overline{U^{\prime}})$. The claim follows. ∎
For each $A$, let $U_{A}$ denote a neighborhood such as the above. By Theorem 3.12 we see that $U_{A}\to\Phi^{\mathrm{KNU}}_{\mathfrak{C}(J)}(U_{A})$ is a proper map of analytic varieties. Let $U_{A}\to W_{A}$ be its Stein factorization. Now define $\Phi_{A}$ to be the restriction of $\overline{\Phi}$ to $U_{A}$. Then $\Phi_{A}:U_{A}\to\overline{\Phi}(U_{A})$ agrees set theoretically with $U_{A}\to W_{A}$. Since the topology on $W_{A}$ is the quotient topology from $U_{A}$, and its structure sheaf $\mathcal{O}^{\mathrm{an}}_{W_{A}}$ agrees with the pushforward of the structure sheaf of $U_{A}$, the Theorem follows. ∎
6. Examples of projective completions in dimension 22
Let us assume $S$ is a quasi-projective surface. The $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow S$ gives the period map $\D\Phi:S\rightarrow\Gamma\backslash$ D. We fix a normal projective completion $\hat{S}$ of $S$ which we do not assume to be a SNC completion. Let $S^{\circ}\subset\hat{S}$ be the complement of codimension-$2$ strata. The main result of this section is: